OpenStudy (anonymous):

Express as a single fraction te exact value of sin75

OpenStudy (anonymous):

not possible

OpenStudy (anonymous):

use a calculator i guess

OpenStudy (anonymous):

I did but it i cant turn the decimal it gives me into a fraction

OpenStudy (anonymous):

oh pellet i got this one

OpenStudy (zepp):

It is possible, using trig identities \(sin(A+B)=\sin A\cos B+\cos B \sin A \)

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

why is pellet gettting typed as pellet ?

OpenStudy (anonymous):

wtf ?

OpenStudy (anonymous):